Hi,
I need some help with these 2 problems:
1. Find three consecutive odd integers such that, the product of the first and the third minus the product of the first and the second, is 11 more than the third number.
2. The annual interest earned by $18,000 is $206 more than the interest earned by $16,000 invested at a rate 0.8% less than the one corresponding to the $18,000. What is the rate of interest for each amount?
Thanks in advance,
Lou

Answer by checkley71(8403)

x(x+4)-x(x+2)=(x+4)-11 or x~2+4x-x~2-2x=x+15 or 4x-2x-x=15 or x=15
proof 15*19-15*17=19+11 or 285-255=30 or 30=30
18000x=206+16000(x-.008) or 18000x=206+16000x-128 or 2000x=206-128 or
2000x=78 or x=.039 or 3.9% for the 18000 investment & 3.9%-.8%=3.1% for the
16000 investment